AM4 -Esercitazione 4 A.A.2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 16 ottobre 2003

Esercizio 1

Calcolare i seguenti integrali indefiniti:

1. ∫ x^{3} arcsin \frac{1}{x} dx

2. ∫ cos(ln x)dx

3. ∫ x^{2} ln \sqrt{1 − x} dx

4. ∫ \frac{x^{2} − 1}{(x − 2)(x^{2} + 1)} dx

5. ∫ \frac{2x − 3}{(x − 2)(x + 1)^{2}} dx

Esercizio 2

Sia f(x, y) = x^{4} + y^{2}. Si consideri il paraboloide

S = {(x, y, z) ∈ ℝ^{3} : 2z = x^{2} + y^{2}, 0 ≤ x^{2} + y^{2} ≤ 1}.

Calcolare ∫_{S} fdσ.

Esercizio 3

Sia I_{n} = ∫ x^{n}e^{−x}dx. Calcolare I_{10}.